write 2 equivalent fraction to 7/14
Answer by sdmmadam@yahoo.com(530)

write 2 equivalent fraction to 7/14
7/14 = 7/(2x7) = 1/2 (cancelling out 7 in the numerator and in the denominator)
And fractions like 3/6, 5/10, 4/8, 8/16, 10/20, etc all on simplification reduce to 1/2 and hence are fractions equivalent to 7/14
Note: any given fraction must be reduced to its most simplified form which is to say all factors common to numerator and denominator should be cancelled out and finally when only 1 is the common factor it is the sign we have reached the most simplified form.